Mike invested $6000 for one year. he invested part of it at 9% and the rest at 11%. At the end of the year he earned $624 in interest. How much did he invest at each rate?
Answer by checkley77(12569)

.11x+.09(6000-x)=624
.11x+540-.09x=624
.02x=624-540
.02x=84
x=84/.02
x=4200 invested @ 11%.
6000-4200=1800 invested @8%.
Proof:
.11*4200+.09*1800=624
462+162=624
624=624